MATH 323-01 INTRODUCTION TO PROBABILITY SPRING 2004 MWF 1:55-2:45 HAYES-HEALY 229

Instructor: François Ledrappier Office: 287 Hurley Office Phone: 631-8331 E-mail: ledrappier.1@nd.edu Textbook: A first course in Probability (6th. edition), by S. Ross Office Hours: Monday 4-5 pm and by appointment.

MATERIAL TO BE COVERED: Most of the textbook.

COURSE WORK:

Homework: Homework problems will be assigned after each class. Assignments will be collected weekly, usually on Wednesdays. Late homework is not accepted. Your lowest homework score for the trimester will be dropped.

Exams: There will be three in-class mid-term tests and a final exam at the following dates:

Test 1: Friday, February 13. Expected covered material: Chapters 1 to 3.

Test 2: Wednesday, March 17. Chapters 4 and 5.

Test 3: Wednesday, April 14. Chapters 6 and 7.

Final: Monday, May 3, 8:00 to 10:00 a.m..

If (and only if) official university activities force you to miss one of these exams, there will be a make-up exam. Please contact the instructor at least one week in advance.

GRADING POLICIES: Homework 45 pts Tests 60 pts each Final 75 pts.